Can you please help me with this problem...Thank you
Suppose Bob owns 2,000 shares of Company X and 10,000 shares of Company Y. The total value of Bob's holdings of these two companies is $340,000.
Suppose Frank owns 8,000 shares of Company X and 6,000 shares of Company Y. The total value of Franks holdings of these two companies is $340,000.
a. Write equations for Bob and Frank's holdings. Use the variables X and Y to represent the values of shares of Company X and Company Y.
b. Solve for the value of a share of Company X and Company Y. Show your work so you can get partial credit even if your final answer is wrong.

Answer by stanbon(75887)

Suppose Bob owns 2,000 shares of Company X and 10,000 shares of Company Y. The total value of Bob's holdings of these two companies is $340,000.
----------------------------------
Bob's holdings: 2000x + 10000y = 340,000 where x and y are the value of one
share of company x or of company y.
------------------------------
Frank's holding: 8000x + 6000y = 340,000 where x and y are the value of
one share of the respective companies.
-----------------------

EQUATION:
2000x + 10000y = 340000
8000x + 6000y = 340000
----------------
Multiply 1st by 4 and subtract from the 2nd to get:
2000x + 10000y = 340000
0x - 34000y = -3*340000
y = $30 per share of company Y
---------------------------------
Substitute to solve for "x":
2000x + 10000*30 = 340000
2000x + 300000 = 340000
x = $20 per share of company X
====================================
Cheers,
Stan H.